ZWITTERIONS AND CAGES: NEW RESULTS FOR NEUTRAL AND PROTONATED WATER CLUSTERS

\underline{JER-LAI KUO}, Department of Chemistry, Ohio State University, Columbus, Ohio 43210; CRIS-TIAN V. CIOBANU, Department of Physics, Ohio State University, Columbus, Ohio 43210; LARS OJAMÄE, Physical Chemistry, Arrhenius Laboratory, Stockholm University, S-106 91 Stockholm, Sweden; ISAIAH SHAVITT and SHERWIN J. SINGER, Department of Chemistry, Ohio State University, Columbus, Ohio 43210.

Commonly held wisdom stipulates that hydrogen bonds between water molecules stabilize a structure by ∼ 5 \frac{kcal}{mol} apiece. Therefore aqueous clusters that differ only by the direction of hydrogen bonds, but otherwise have the same number of Hbonds and placement of oxygen atoms should have roughly the same energy, and spread of energies should be far less than ∼ 5 \frac{kcal}{mol} per water molecule. This belief lies behind calulations performed to date for the (H_{2}O)_{20} dodecahedron, and the associated H^{+}(H_{2}O)_{21} formed by adding a hydronium ion, in which only one, or a handful of arbitrarily chosen hydrogen bond arrangements were selected for study. Presumably the H-bond arrangement in the dodecahedral cage has minor effect on the chemistry of this cluster. In this work we show that H-bond topology strongly affects the structure and energy of (H_{2}O)_{20}. This was implicated in our previous studies of (H_{2}O)_{20} ^{a} ^{b} using empirical and semi-empirical models, and is now confirmed by ab initio electronic structure calculations. Furthermore, we find that the H-bond arrangement strongly affects not just the structure and energy of (H_{2}O)_{20}, but also its chemistry. In fact, a seeming innocuous rearrangement of the H-bonds in (H_{2}O)_{20} leads to spontaneous autoionization of this structure, producing spatially separated hydrogen and hydroxide ions in the cluster. Thermal behavior and zwitterion formation in smaller aqueous clusters is also considered.

aShannon McDonald, Lars Ojamäe and Sherwin J. Singer, J. Phys. Chem. A 102, 2824 (1998)

bJer-Lai Kuo, James V. Coe, Sherwin J. Singer, Yehuda B. Band, and Lars Ojamäe, J. Chem. Phys. 114, 2527 (2001)